OpenStudy (stormoshawty):

Solve 3x-9y=3 6x-3y=-24 using substitution.

OpenStudy (shamil98):

3x - 9y = 3 6x - 3y = 24 We need to find x or y first. start with substituting with x. 3x = 9y + 3 x = 3y + 1 substitute 6(3y+1) - 3y = 24 solve for y now

OpenStudy (stormoshawty):

I got \[y = \frac{ 6 }{ 5 }\]

OpenStudy (shamil98):

18y + 6 - 3y = 24 15y = 18 y = 6/5 yep correct

OpenStudy (shamil98):

3x - 9y = 3 substitute back in to solve for x. 3x - 9 (6/5) = 3

OpenStudy (stormoshawty):

\[y=\frac{ 23 }{ 5 }?\]

OpenStudy (stormoshawty):

Oops, i meant x.

OpenStudy (shamil98):

3x - 54/5 = 3 3x = 3 + 54/5 3x = 69/5 x = 23/5 yep good work!

OpenStudy (stormoshawty):

Thanks, I have like 2 more.

OpenStudy (shamil98):

yeah, just post it seperately